Calculate the return or yield on an deposit of  $3000 that grows to $3650 compounding annually for 5 years.
Answer #1

Here present value = 3000$
Future value = FV = 3650$
N = number of year = 5 years
r = rate of interest = ?
FV = PV(1+r)^n
=3650 = 3000(1+r)^5
=1.216667 = (1+r)^5

=1.216667(15) = (1 + r) 52(1/5)


=1.216667 ^ 0.2 = 1+r
=1.040002=1+r
r = 0.040002
i.e 4%